Find the value of a2 for the recurrence relation an=17an-1+30n, where a0=3



Find the value of a3 for the recurrence relation an=17an-1+30n, where a0=3



Find the value of a1 for the recurrence relation an=17an-1+30n, where a0=3



What would be the hypothesis of the mathematical induction for x(x + 1) < x! , where x ≥ 7?



If P(k) = k2(k + 2)(k – 1) is true, then what is P (k + 1)?

1
Expert's answer
2022-06-09T14:37:57-0400

1.


"a_1=17(3)+30(1)=81"

"a_2=17(81)+30(2)=1437"

2.


"a_3=17(1437)+30(3)=24519"

3.


"a_1=17(3)+30(1)=81"

4.

It is assumed that at "x = k," "P(k)" holds


"k(k+1)<k!"



5.

If "P(k) = k^2(k + 2)(k \u2013 1)" is true, then it must be shown that "P(k + 1)" is true, namely, that

"(k+1)^2(k+1 + 2)(k+1 \u2013 1)"

is also true.

Simplify


"k(k+1)^2(k+3)"

is also true.